# When Is Degree Enough? Bounds on Degree-Eigenvector Misalignment in Assortative Structured Networks

### v1

@@ Abstract
A tight alignment between the degree vector and the leading eigenvector arises naturally in networks with neutral degree mixing and the absence of local structures. Many real-world networks, however, violate both conditions. We derive bounds on the divergence between the degree vector and the eigenvector in networks with degree assortativity and local mesoscopic structures such as communities, core-peripheries, and cycles. Our approach is constructive. We design sufficiently general degree-preserving rewiring algorithms that start from a neutral benchmark and monotonically increase assortativity and the strength of local structures, with each step inducing a perturbation of the adjacency matrix. Using the Stewart–Sun Perturbation Bound, together with explicit spectral-norm control of the rewiring steps, we derive upper bounds on the angle between the eigenvector and the degree vector for modest levels of assortativity and local structures. Our analytical bounds delineate regions of ‘spectral safety’ in which a node’s degree can be used as a reliable measure of its systemic importance in real-world networks.
@@ 1 Introduction
Over the last quarter century, many disciplines—including the social and biological sciences—have embraced the idea that a node’s systemic importance depends not only on how many neighbors it has, but also on where it sits in the network as a whole. In such settings, eigenvector-based measures of centrality provide a natural notion of influence or vulnerability. In practice, however, computing eigenvector centrality requires essentially complete knowledge of the network, which is rarely available. By contrast, degree information is often observed or can be estimated reasonably well. Which is why economists, epidemiologists, and other applied researchers frequently use the degree vector as a proxy for the leading eigenvector. This is entirely reasonable when the two are close. Unfortunately, however, the conditions under which degree and eigenvector coincide are quite restrictive. The network must not only exhibit no assortativity, it must also be devoid of all local structures, including communities, core-periphery patterns, and directed cycles. One would be hard-pressed to find a real-world network that meets these conditions. Firm buyer–seller networks, for instance, exhibit a clear core-periphery structure. Networks of disease transmission sit inside rich social graphs with intricate communities and friendship cycles [^1] See Salathé and Jones (2010) and Volz et al. (2011) for evidence on local structures within the social networks of epidemiological transmission. See Chakraborty et al. (2018) and Bacilieri et al. (2025) for empirical evidence on assortativity and local structures in firm buyer-seller networks.. In such circumstances, the leading eigenvector can diverge from the degree profile, with the degree vector becoming a progressively worse proxy as the network departs from the neutral benchmark. Ideally, we would like to know how large this divergence is before using the degree vector as a proxy for the eigenvector. Which is precisely the question we address in this paper. We derive explicit bounds on the divergence between the eigenvector and the degree vector when the network departs from the neutral benchmark by modest amounts of assortativity, community structure, core-periphery pattern, and directed cycles.
@@
Nearly all of what is said in this paper is built upon the Stewart-Sun Perturbation Bound, which tells us how much the eigenvector can move under a perturbation of the adjacency matrix. In its simplest form, the theorem states that if $\mathbf{A}$ is diagonalizable with a simple leading eigenvalue and we form a perturbed matrix $\widetilde{\mathbf{A}}:=\mathbf{A}+\boldsymbol{\Omega}$, then the sine of the angle between the corresponding leading eigenvectors is bounded above by a constant times $\|\boldsymbol{\Omega}\|_{2}$, the spectral norm of the perturbation. Taken at face value, this is a statement about eigenvectors before and after perturbation. It says nothing about degree vectors. Note, however, that under the neutral benchmark, the leading eigenvector of $\mathbf{A}$ is well approximated by the degree vector. Therefore, if we restrict perturbations to degree-preserving rewirings, the Stewart-Sun bound on the distance between the eigenvectors of $\mathbf{A}$ and $\widetilde{\mathbf{A}}$ is equivalent to a bound on the distance between the degree vector and the eigenvector of the perturbed graph $\widetilde{\mathbf{A}}$. We develop this insight by constructing four families of degree-preserving rewiring schemes that progressively introduce (i) assortative mixing, (ii) community structures, (iii) core-periphery patterns, and (iv) directed cycles. For each scheme we explicitly bound $\|\boldsymbol{\Omega}\|_{2}$ in terms of standard scalar measures of these features—Newman’s assortativity coefficient for degree correlations, a modularity-type quantity for communities, a core-periphery contrast parameter, and normalized counts of cycles. Substituting these into the Stewart–Sun inequality yields explicit, interpretable bounds on the divergence between degree and eigenvector centrality as a function of the strength of assortativity and local structure. These are the first analytical results that link commonly used network statistics to quantitative guarantees on the accuracy of degree-based proxies for eigenvector centrality.
@@
The paper is organized as follows. Section 2 sets notation and collects the spectral preliminaries needed for our main arguments, including the Stewart–Sun perturbation bound. Section 3 bounds the moments of the angle of deviation with the moments of network statistics. Section 4 details the assortativity-inducing degree-preserving rewiring process. It then applies the moment-bound proposition to assortativity, while noting how the constants in the bound depend on certain structural features of the network. Section 5 details the degree-preserving rewiring processes used to generate three distinct local structures: communities, core-periphery patterns, and cycles. We generalize each of these structures, with $K$ communities, $k$ cycles, and perhaps most interestingly, fractal core-periphery wherein the periphery itself consists of ‘miniature’ core-periphery structures. We then apply the moment bound proposition to bound the angle of deviation with measures of communities, core-periphery, and cycles. Section 6 discusses how the heavy tails of the degree distribution shape the bounds on the angle of deviation between the degree-vector and eigenvector. Section 7 offers concluding remarks.
@@ 2.2 Neutral matrix and its alignment with the degree vector
A baseline neutral network is an unweighted directed graph on $n$ nodes with adjacency matrix $\mathbf{A}\in\mathbb{R}_{\geq 0}^{n\times n}$. We assume that $\mathbf{A}$ is irreducible, that the Perron root $\lambda_{1}(\mathbf{A})$ is simple, and that $\mathbf{A}$ is diagonalizable, so that the Perron eigenvectors are well-defined up to scale. [^2] We work throughout in the Perron–Frobenius regime and assume that the rewiring steps we consider are sufficiently few and local that the graph remains irreducible along the trajectory. Equivalently, one may restrict attention to rewiring moves that preserve strong connectivity (e.g., by rejecting swaps that break irreducibility).
@@ 2.3 Perturbation matrices and their spectral norms
$\boldsymbol{\Omega}^{(t)}\;:=\;\sum_{i=1}^{t}\boldsymbol{\Delta}^{(i)}.$
@@
$\mathbf{A}^{(t)}\;:=\;\mathbf{A}\;+\;\boldsymbol{\Omega}^{(t)}.$
@@ 2.4 Stewart–Sun eigenvector perturbation bound
When the underlying matrix is clear from context, we write $\kappa$ for $\kappa(\mathbf{A})$. This quantity measures how far the eigenvector matrix $\mathbf{V}$ is from being orthogonal [^3] For directed networks, adjacency matrices need not admit an orthogonal eigenbasis. In particular, for a $\{0,1\}$ adjacency matrix $\mathbf{A}$ one has $(\mathbf{A}^{\top}\mathbf{A})_{ii}=d_{i}^{\mathrm{in}}$ and $(\mathbf{A}\mathbf{A}^{\top})_{ii}=d_{i}^{\mathrm{out}}$. Thus if $\mathbf{A}$ were (real) normal and hence orthogonally diagonalizable, $\mathbf{A}^{\top}\mathbf{A}=\mathbf{A}\mathbf{A}^{\top}$ would force $d_{i}^{\mathrm{in}}=d_{i}^{\mathrm{out}}$ for every node $i$. Therefore any node-level divergence $d_{i}^{\mathrm{in}}\neq d_{i}^{\mathrm{out}}$ rules out $\kappa(\mathbf{A})=1$ (though degree-balance alone does not guarantee $\kappa(\mathbf{A})=1$). : in particular, $\kappa(\mathbf{A})=1$ when $\mathbf{V}$ is orthogonal, i.e.,
@@ 3 Bounding moments of angle of deviation with moments of network statistic
The interplay of (i) and (ii) can therefore produce a trajectory for $\theta_{t}$ that wanders within $0,\pi/2]$ despite the strict monotonicity of $\phi(\mathbf{A}^{(t)})$ along each step [^6] This conjecture helps explain why the empirical evidence on the matter is mixed. One line of work shows that as local or mesoscopic structure accumulates, the leading eigenvector may concentrate on a small set of vertices and drift away from degree-based rankings (Sharkey, 2019; Pastor-Satorras and Castellano, 2016). But the opposite tendency can also occur: within a fixed degree sequence, there are neutral realizations in which degree and eigenvector centrality remain close. This includes maximum-entropy/BFD graphs (Atay and Bıyıkoğlu, forthcoming). Some rewiring protocols effectively add structure in ways that track degree rather than compete with it. In short, different networks and protocols can therefore generate different outcomes. Our conjecture gives reasons for why the empirical record does not point in a single direction.
@@
$\mathbb{E}\bigl[\theta(\mathbf{A}^{(t+1)})\mid\mathbf{A}^{(t)}\bigr]\ \geq\ \theta(\mathbf{A}^{(t)}),\qquad\text{with }\mathbf{A}^{(t+1)}\in\widetilde{\mathcal{B}}^{(t)}t$
@@
(i) Linear spectral sensitivity (Stewart–Sun). There is a constant $L_{\theta}\lesssim\kappa_{*}/\gamma_{*}$ such that for all $A\in S_{0}$,
@@
(ii) Participation-controlled perturbation size (no explicit $n$). Under $r_{\phi}$-bounded participation, the cumulative perturbation $\Omega^{(t)}=\sum_{s=0}^{t-1}\Delta^{(s)}$ satisfies
@@ 4 Assortativity
where $C_{p,q}$ is chosen large enough to absorb normalization conventions and any residual slack terms from the affine comparison.
If the in/out maxima are comparable and the common degree tail is power-law with exponent $\alpha>1$ (so $d_{\max}\asymp n^{1/\alpha}$ up to slowly varying factors), then $\Lambda_{p,q}\asymp\nu_{p,q}d_{\max}^{2}$ and hence
$M_{p,q}\;\lesssim\;\widetilde{C}_{p,q}\,\nu_{p,q}\,\frac{\kappa_{*}}{\gamma_{*}}\,r\,n^{2/\alpha}$
so heavier tails (smaller $\alpha$) enlarge the time-uniform envelope for $\mathbb{E}[\theta_{t}]$ through increased extremal-degree leverage. $\circ$
@@ 5 Local structures
We now turn to local and mesoscopic structures that can pull the leading eigenvector away from the degree proxy. We focus on three canonical patterns: communities (modularity), core-periphery organization, and cycles. For each pattern we specify a degree-preserving rewiring rule that amplifies the structure, express the resulting change as a cumulative perturbation of the adjacency matrix, and then invoke the Stewart-Sun perturbation bound (Result 1), together with our operator-norm control of the perturbation, to bound the induced degree-eigenvector deviation.
@@ 5.1 Communities
If the in/out maxima are comparable and the common degree tail is power-law with exponent $\alpha>1$ (so $d_{\max}^{\mathrm{out}}\asymp d_{\max}^{\mathrm{in}}\asymp d_{\max}\asymp n^{1/\alpha}$ up to slowly varying factors), then $\Lambda_{\mathrm{com}}\asymp C_{\mathrm{com}}(\mathcal{C})\,d_{\max}^{2}$ and hence
$M_{\mathrm{com}}\;\lesssim\;\widehat{C}_{\mathrm{com}}(\mathcal{C})\,\frac{\kappa_{*}}{\gamma_{*}}\,r_{\mathrm{com}}\,n^{2/\alpha},$
so heavier tails (smaller $\alpha$) enlarge the time-uniform envelope for $\mathbb{E}[\theta_{t}]$ through increased extremal-degree leverage under degree-preserving swaps. $\circ$
@@ 5.2 Core-periphery
If the in/out maxima are comparable and the common degree tail is power-law with exponent $\alpha>1$ (so $d_{\max}^{\mathrm{out}}\asymp d_{\max}^{\mathrm{in}}\asymp d_{\max}\asymp n^{1/\alpha}$ up to slowly varying factors), then $\Lambda_{\mathrm{cp}}\asymp C_{\mathrm{cp}}(\mathcal{H},\mathcal{L})\,d_{\max}^{2}$ and hence
$M_{\mathrm{cp}}\;\lesssim\;\widehat{C}_{\mathrm{cp}}(\mathcal{H},\mathcal{L})\,\frac{\kappa_{*}}{\gamma_{*}}\,r_{\mathrm{cp}}\,n^{2/\alpha}$
so heavier tails (smaller $\alpha$) enlarge the time-uniform envelope for $\mathbb{E}[\theta_{t}]$ through increased extremal-degree leverage under degree-preserving swaps. $\circ$
@@ 5.3 Cycles
If the in/out maxima are comparable and the common degree tail is power-law with exponent $\alpha>1$ (so $d_{\max}^{\mathrm{out}}\asymp d_{\max}^{\mathrm{in}}\asymp d_{\max}\asymp n^{1/\alpha}$ up to slowly varying factors), then $\Lambda_{k\text{-}\mathrm{cyc}}\asymp C_{k}d_{\max}^{2}$ and hence
$M_{k\text{-}\mathrm{cyc}}\;\lesssim\;\widehat{C}_{k}\,\frac{\kappa_{*}}{\gamma_{*}}\,r_{k\text{-}\mathrm{cyc}}\,n^{2/\alpha}$
so heavier tails (smaller $\alpha$) enlarge the time-uniform envelope for $\mathbb{E}[\theta_{t}]$ by increasing the combinatorial leverage available for closing cycles around hubs. $\circ$
6 Heavy-tailed degrees and eigenvector localization
@@
A subtle but practically important point is that the constants in our moment bounds are not universal, and this is exactly where the four statistics truly part ways. All four bounds share the same backbone: (i) eigenvector responsiveness enters through distortion and spectral-gap parameters $(\kappa,\gamma)$, and (ii) the cumulative size of the perturbation is controlled by the bounded-participation operator norm $r$. Where they differ is in the statistic-to-angle sensitivity—how changes in the chosen global statistic $\phi$ translate into bounds on the deviation angle $\theta(\mathbf{A})$. The algebraic form of the moment transfer is stable, but its scale is dictated by how sensitive a statistic is under degree-preserving swaps and, conversely, how much eigenvector rotation those swaps can actually induce. A convenient proxy for this sensitivity is given by the affine parameters, whose exact functional form and magnitude is inherently statistic-specific.
Put differently, there are really two layers at work. The first layer is purely spectral: $\kappa/\gamma$ measures how fragile the leading eigenvector is to perturbations of a given operator norm. The second layer is combinatorial/geometric: it tells you how much operator-norm perturbation (and hence how many degree-preserving edits, under the $r$-bounded participation rule) is needed to achieve a prescribed change in $\phi$. The statistics differ primarily in this second layer—some can be moved substantially by rearranging a small number of edges concentrated around a few vertices, while others only shift appreciably after many more, and more dispersed, changes.
Assortativity.

### v2

@@ Abstract
A tight alignment between the degree vector and the leading eigenvector arises naturally in networks with neutral degree mixing and the absence of local structures. Many real-world networks, however, violate both conditions. We derive bounds on the divergence between the degree vector and the eigenvector in networks with degree assortativity and local mesoscopic structures such as communities, core-peripheries, and cycles. Our approach is constructive. We design sufficiently general degree-preserving rewiring algorithms which begins from a neutral benchmark to monotonically change assortativity and the strength of local structures, with each step inducing a perturbation of the adjacency matrix. Using the Stewart–Sun Perturbation Bound, together with explicit spectral-norm control of the rewiring steps, we derive upper bounds on the angle between the eigenvector and the degree vector for modest levels of assortativity and local structures. Our analytical bounds delineate regions of ‘spectral safety’ in which a node’s degree can be used as a reliable measure of its systemic importance in real-world networks. We also substantiate our analytical bounds with numerical simulations that compute the exact angles of deviation.
@@ 1 Introduction
Over the last quarter century, many disciplines—including the social and biological sciences—have embraced the idea that a node’s systemic importance depends not only on how many neighbors it has, but also on where it sits in the network as a whole. In such settings, eigenvector-based measures of centrality provide a natural notion of influence or vulnerability. In practice, however, computing eigenvector centrality requires essentially complete knowledge of the network, which is rarely available. By contrast, degree information is often observed or can be estimated reasonably well. Which is why applied researchers sometimes use the degree vector as a proxy for the leading eigenvector. This is entirely reasonable when the two are close. Unfortunately, however, the conditions under which degree and eigenvector coincide are quite restrictive. The network must not only exhibit no assortativity, it must also be devoid of all local structures, including communities, core-periphery patterns, and directed cycles. One would be hard-pressed to find a real-world network that meets these conditions. Firm buyer–seller networks, for instance, exhibit a clear core-periphery structure. Networks of disease transmission sit inside rich social graphs with intricate communities and friendship cycles [^1] See Salathé and Jones (2010) and Volz et al. (2011) for evidence on local structures within the social networks of epidemiological transmission. See Chakraborty et al. (2018) and Bacilieri et al. (2025) for empirical evidence on assortativity and local structures in firm buyer-seller networks.. In such circumstances, the leading eigenvector can diverge from the degree profile, with the degree vector becoming a progressively worse proxy as the network departs from the neutral benchmark. Ideally, we would like to know how large this divergence is before using the degree vector as a proxy for the eigenvector. Which is precisely the question we address in this paper. We derive explicit bounds on the divergence between the eigenvector and the degree vector when the network departs from the neutral benchmark by modest amounts of assortativity, community structure, core-periphery pattern, and directed cycles. While our primary focus is on unweighted graphs, we also demonstrate how this framework naturally generalizes to weighted directed networks.
@@
Nearly all of what is said in this paper is built upon the Stewart-Sun Perturbation Theorem, which tells us how much the eigenvector can move under a perturbation of the adjacency matrix. In its simplest form, the theorem states that if $\mathbf{A}$ is diagonalizable with a simple leading eigenvalue and we form a perturbed matrix $\widetilde{\mathbf{A}}:=\mathbf{A}+\boldsymbol{\Omega}$, then the sine of the angle between the corresponding leading eigenvectors is bounded above by a constant times $\|\boldsymbol{\Omega}\|_{2}$, the spectral norm of the perturbation. Taken at face value, this is a statement about eigenvectors before and after perturbation. It says nothing about degree vectors. Note, however, that under the neutral benchmark, the leading eigenvector of $\mathbf{A}$ is well approximated by the degree vector. Therefore, if we restrict perturbations to degree-preserving rewirings, the Stewart-Sun bound on the distance between the eigenvectors of $\mathbf{A}$ and $\widetilde{\mathbf{A}}$ is equivalent to a bound on the distance between the degree vector and the eigenvector of the perturbed graph $\widetilde{\mathbf{A}}$. We develop this insight by constructing four families of degree-preserving rewiring schemes that progressively introduce (i) assortative mixing, (ii) community structures, (iii) core-periphery patterns, and (iv) directed cycles. For each scheme we explicitly bound $\|\boldsymbol{\Omega}\|_{2}$ in terms of standard scalar measures of these features—Newman’s assortativity coefficient for degree correlations, a modularity-type quantity for communities, a core-periphery contrast parameter, and normalized counts of cycles. Substituting these into the Stewart–Sun inequality yields explicit, interpretable bounds on the divergence between degree and eigenvector centrality as a function of the strength of assortativity and local structure. These are the first analytical results that link commonly used network statistics to quantitative guarantees on the accuracy of degree-based proxies for eigenvector centrality.
@@
The paper is organized as follows. Section 2 sets notation and collects the spectral preliminaries needed for our main arguments, including the Stewart–Sun perturbation bound. Section 3 bounds the moments of the angle of deviation with the moments of network statistics. Section 4 details the assortativity-inducing degree-preserving rewiring process. It then applies the moment-bound proposition to assortativity, while noting how the constants in the bound depend on certain structural features of the network. Section 5 details the degree-preserving rewiring processes used to generate three distinct local structures: communities, core-periphery patterns, and cycles. We generalize each of these structures, with $K$ communities, $k$ cycles, and perhaps most interestingly, fractal core-periphery wherein the periphery itself consists of ‘miniature’ core-periphery structures. We then apply the moment bound proposition to bound the angle of deviation with measures of communities, core-periphery, and cycles. Section 6 discusses how the heavy tails of the degree distribution shape the bounds on the angle of deviation between the degree-vector and eigenvector. Section 7 offers concluding remarks. Appendix A extends the analytical results to weighted directed graphs by considering the continuous analogue of our discrete rewiring setting. Appendix B presents numerical simulations that estimate the exact angle of deviation for specific quantums of assortativity and community structures. These simulations support our analytical results about upper bounds on the angle of deviation and further clarify the mechanisms involved in generating these deviations.
@@ 2.2 Neutral matrix and its alignment with the degree vector
A baseline neutral network is an unweighted directed graph on $n$ nodes with adjacency matrix $\mathbf{A}\in\mathbb{R}_{\geq 0}^{n\times n}$. We assume that $\mathbf{A}$ is irreducible, that the Perron root $\lambda_{1}(\mathbf{A})$ is simple, and that $\mathbf{A}$ is diagonalizable, so that the Perron eigenvectors are well-defined up to scale. [^2] We throughout assume that the rewiring steps we consider are sufficiently few and local that the graph remains irreducible along the trajectory. Equivalently, one may restrict attention to rewiring moves that preserve strong connectivity (e.g., by rejecting swaps that break irreducibility).
@@ 2.3 Perturbation matrices and their spectral norms
$\boldsymbol{\Omega}^{(t)}\;:=\;\sum_{i=1}^{t}\boldsymbol{\Delta}^{(i)}$
@@
$\mathbf{A}^{(t)}\;:=\;\mathbf{A}\;+\;\boldsymbol{\Omega}^{(t)}$
@@ 2.4 Stewart–Sun eigenvector perturbation bound
When the underlying matrix is clear from context, we write $\kappa$ for $\kappa(\mathbf{A})$. This quantity measures how far the eigenvector matrix $\mathbf{V}$ is from being orthogonal [^3] For directed networks, adjacency matrices need not admit an orthogonal eigenbasis. Here “normal” means that $\mathbf{A}^{\top}\mathbf{A}=\mathbf{A}\mathbf{A}^{\top}$, which is the relevant benchmark for $\kappa(\mathbf{A})=1$. For a $\{0,1\}$ adjacency matrix $\mathbf{A}$, one has $(\mathbf{A}^{\top}\mathbf{A})_{ii}=d_{i}^{\mathrm{in}}$ and $(\mathbf{A}\mathbf{A}^{\top})_{ii}=d_{i}^{\mathrm{out}}$. Hence normality implies $d_{i}^{\mathrm{in}}=d_{i}^{\mathrm{out}}$ for every node $i$. Therefore any node-level divergence $d_{i}^{\mathrm{in}}\neq d_{i}^{\mathrm{out}}$ rules out $\kappa(\mathbf{A})=1$, although degree balance alone does not guarantee normality.
: in particular, $\kappa(\mathbf{A})=1$ when $\mathbf{V}$ is orthogonal, i.e.,
@@ 3 Bounding moments of angle of deviation with moments of network statistic
The interplay of (i) and (ii) can therefore produce a trajectory for $\theta_{t}$ that wanders within $0,\pi]$ despite the strict monotonicity of $\phi(\mathbf{A}^{(t)})$ along each step [^6] This conjecture helps explain why the empirical evidence on the matter is mixed. One line of work shows that as local or mesoscopic structure accumulates, the leading eigenvector may concentrate on a small set of vertices and drift away from degree-based rankings (Sharkey, 2019; Pastor-Satorras and Castellano, 2016). But the opposite tendency can also occur: within a fixed degree sequence, there are neutral realizations in which degree and eigenvector centrality remain close. This includes maximum-entropy/BFD graphs (Atay and Bıyıkoğlu, forthcoming). Some rewiring protocols effectively add structure in ways that track degree rather than compete with it. In short, different networks and protocols can therefore generate different outcomes. Our conjecture gives reasons for why the empirical record does not point in a single direction.
@@
$\mathbb{E}\bigl[\theta(\mathbf{A}^{(t+1)})\mid\mathbf{A}^{(t)}\bigr]\ \geq\ \theta(\mathbf{A}^{(t)}),\qquad\text{with }\mathbf{A}^{(t+1)}\in\widetilde{\mathcal{B}}^{(t)}$
@@
(i) Linear spectral sensitivity. There is a constant $L_{\theta}\lesssim\kappa_{*}/\gamma_{*}$ such that for all $A\in S_{0}$,
@@
(ii) Participation-controlled perturbation size. Under $r_{\phi}$-bounded participation, the cumulative perturbation $\Omega^{(t)}=\sum_{s=0}^{t-1}\Delta^{(s)}$ satisfies
@@ 4 Assortativity
where $C_{p,q}$ is chosen large enough to absorb normalization conventions and any residual slack terms from the affine comparison. $\circ$
@@ 5 Local structures
We now turn to local and mesoscopic structures that can pull the leading eigenvector away from the degree proxy. We focus on three canonical patterns: communities (modularity), core-periphery organization, and cycles. For each pattern we specify a degree-preserving rewiring rule that amplifies the structure, express the resulting change as a cumulative perturbation of the adjacency matrix, and then invoke the Stewart-Sun Perturbation Theorem (Result 1), together with our operator-norm control of the perturbation, to bound the induced degree-eigenvector deviation.
@@ 5.1 Communities
$\circ$
@@ 5.2 Core-periphery
$\circ$
@@ 5.3 Cycles
$\circ$
6 Power-law tails and statistic-specific bounds
Many large real-world networks exhibit pronounced degree heterogeneity, often well approximated in their upper tails by power laws. For our purposes, the importance of this empirical regularity is not merely descriptive. It identifies the regime in which a small number of vertices carry a disproportionate share of edge-ends, and hence the regime in which degree-preserving rewiring can have especially uneven and high-leverage effects on global network statistics. This is also the regime in which the constants appearing in our moment bounds become most informative.
@@ 6 Power-law tails and statistic-specific bounds
A practically important point is that these constants are not universal. All four statistics we study share the same spectral backbone: eigenvector responsiveness enters through the conditioning and separation parameters $(\kappa,\gamma)$, while the cumulative size of the rewiring enters through the bounded-participation operator-norm budget $r$. Where the statistics differ is in how a change in the chosen global statistic $\phi$ translates into a bound on the deviation angle $\theta(\mathbf{A})$. The ratio $\kappa/\gamma$ captures the common spectral part of this relation by measuring how fragile the Perron eigenvector is to perturbations of a given operator norm. What varies across statistics is the combinatorial part: how much such perturbation, and hence how many degree-preserving edits under the $r$-bounded participation rule, are needed to move the statistic by a given amount. This difference is reflected in the affine constants appearing in the moment bounds. Their formal role is the same across statistics, but their magnitude is not, since it depends on how sensitive the statistic is to degree-preserving swaps and on how much eigenvector rotation those swaps can induce. In the power-law regime, this statistic-specific component becomes especially transparent, because the dominant contribution comes from extremal degrees. If
$d_{\max}^{\mathrm{out}}\asymp d_{\max}^{\mathrm{in}}\asymp d_{\max}\asymp n^{1/\alpha},\qquad\alpha>1,$
up to slowly varying factors, then each statistic acquires a common heavy-tail scaling of order $d_{\max}^{2}\asymp n^{2/\alpha}$, with the remaining constant determined by the particular feature being rewired. We now examine how this works for assortativity, community structure, core–periphery structure, and cycles.
Assortativity